# MELA: Multilingual Evaluation of Linguistic Acceptability

### v2

@@ 4 Evaluating LLMs with MELA
In this section, we report the performance of several LLMs, open-sourced or close-sourced, on MELA.
@@ 4.1 Experimental Settings
For open-sourced models, we consider BLOOMZ (Scao et al., 2022; Muennighoff et al., 2023), two instruction finetuned variants of mT5 (Xue et al., 2021)—namely mTk (Wang et al., 2022) and mT0 (Muennighoff et al., 2023)—and Baichuan2-Chat (Yang et al., 2023). BLOOMZ is both pretrained and finetuned on 46 languages, which only covers 5 languages in MELA: English, Chinese, French, Spanish, and Arabic222Muennighoff et al. (2023) examine BLOOM’s pretraining corpus ROOTS and estimate it to also contain a small amount of Russian, German, Italian, and Japanese.. The pretraining corpus of mT5 includes all 10 languages in MELA, but mT0 is finetuned on the same instruction dataset as BLOOMZ. mTk’s finetuning data, on the other hand, covers nine languages in MELA (the left-out one is Icelandic) and includes the English CoLA dataset. For Baichuan2, the exact language distribution of pretraining and finetuning data is not disclosed. For close-sourced models, we consider GPT-3.5 and GPT-4 (OpenAI, 2023).
When evaluating mTk, we use 2-shot prompts following the format of its finetuning dataset. For other models, we consider both 0-shot and 2-shot evaluation. More details about the prompts used for evaluating these models are given in Appendix A.
@@ 4.2 Results
The results of LLMs’ performance on MELA are given in Table 3. We make the following observations.
On average, GPT-4 performs on par with supervised models.
It performs only five points below XLM-R in the zero-shot setting, and only one point below it in two-shot setting. On German, French and Spanish even the zero-shot performance of GPT4 is noticeably higher than XLM-R. On Arabic and Icelandic, however, it lags behind even in the two-shot setting, suggesting that GPT-4 may be weaker at understanding these languages.
In few-shot evaluation, using only English examples hurts performance.
As indicated by the results of GPT-3.5, GPT-4 and Baichuan-2, prompting with two English examples leads to even lower performance than 0-shot evaluation. In contrast, prompting with English instructions and in-language examples boosts performance. This suggests that these LLMs fail to transfer the concept of linguistic acceptability acquired from the in-context examples across languages.
Instruction finetuning on acceptability judgements helps cross-lingual transfer.
Of the open-source instruction-finetuned models, mTk performs much better than other models, as its finetuning dataset includes English CoLA. However, mTk also performs much better in non-English examples, and its performance gap between prompting with in-language and English examples is much smaller compared with Baichuan or GPT, suggesting that finetuning on acceptability judgements may unlock the ability of cross-lingual generalization in this task.
5 Cross-lingual Transfer of Linguistic Acceptability
In this section, we investigate cross-lingual transfer in linguistic acceptability by finetuning XLM-RoBERTa (Conneau et al., 2020), which is a multilingual version of RoBERTa (Liu et al., 2019) pretrained on 2.5TB CommonCrawl corpus covering one hundred languages.
@@ 5.1 Experimental Settings
To observe the transfer of acceptability judgements across languages, we train the model on one language, and evaluate it on all ten languages’ development sets. Further training details can be found in Appendix B. We report the median MCC of seven runs for all results to mitigate inter-run variance.
@@ 5.2 Results
The main results of cross-lingual transfer in acceptability judgements are presented in Table 4. Here we make several key observations.
The first is that the ability to perform judgement of linguistic acceptability can be transferred non-trivially across languages, as indicated by the last column of Table 4. The second is that in-language training significantly boosts XLM-R’s performance. Comparing the figures on the diagonal with the last row, this is most prominent for the four high-resource languages. For example, when evaluating on English, training on English leads to 71.66 MCC, compared with an average of 31.60 when training on other nine languages. For low-resource languages, the gap is smaller, but still notable (e.g. for Icelandic the comparison is 25.12 against 17.13). However, we note that for Spanish and Japanese, the highest performance is not obtained when training in-language, but training on English. This leads to our third observation—the number of training samples matters. As indicated by the antepenultimate and penultimate lines of Table 4, when training on high-resource languages, XLM-R obtains an average of 29.69 MCC, compared with 17.18 when training on low-resource languages.
6 Multilingual Fine-tuning with Linguistic Acceptability
In the previous two sections, we investigated the transfer of linguistic acceptability with both LLMs and supervised XLM-R, and found that in-language training data or in-context examples play a key role in learning linguistic acceptability. To further assess the importance of in-language data, we experiment with training on multiple languages. To compensate the impact of training set size, we first downsample data in all languages to the same amount, and then finetune XLM-R on different combinations of languages. Training details are provided in Appendix B.
6.1 Experimental Setting
We downsample sentences in each language to the same number, and train XLM-R in three settings: 1) in-language finetuning; 2) all-language finetuning, where the model is trained on a mixture of data containing an equal number of sentences from ten languages; and 3) all-but-in-language finetuning, where the model is trained on a mixture of data containing an equal number of sentences from nine languages, except the one being evaluated on. Additional experiments on bilingual training are provided in Appendix D.1.
6.2 Results
The results on ten languages’ validation sets are plotted in Figure 1. When trained and evaluated on the same language, the model’s performance scales smoothly with the number of training samples. When trained on sentences from other nine languages, however, the performance saturates at around 500-1000 training samples, consistent with previous findings about multi-task finetuning (Wang et al., 2022). When trained on all ten languages, the performance scales more steadily than all-but-in-language training, but still lags behind in-language training by a large margin, indicating the importance of in-language training data.
7 Edge Probing
In this section, we experiment on several probing tasks to explore whether training on acceptability judgment tasks improves syntax-related abilities of the pre-trained XLM-R.
Task base en it ru zh pos 92.87 93.77 93.47 93.17 93.95 dep 89.41 90.34 90.13 89.92 89.86 const 78.54 79.10 78.44 79.26 78.96 name 93.49 94.23 93.34 94.53 94.08 srl 77.93 82.34 80.00 81.24 80.28 coref 83.84 85.55 84.12 83.98 84.53 avg 86.01 87.56 86.58 87.02 86.94
Probing task Part-of-speech tagging Depedency labeling eval / train en it ru zh avg en it ru zh avg en base 92.87 75.77 65.63 43.33 69.40 89.41 74.99 60.67 40.05 66.28 XLM-R 93.77 81.43 68.22 44.66 72.02 90.34 77.40 61.84 45.44 68.76 it base 83.26 94.61 66.90 38.73 70.88 78.17 91.50 60.65 32.35 65.67 XLM-R 85.6 95.71 63.73 39.70 71.19 83.56 92.46 62.85 37.31 69.05 ru base 82.97 79.90 95.53 53.18 77.90 77.72 78.86 90.90 42.77 72.56 XLM-R 85.42 81.01 95.43 54.06 78.98 80.65 81.27 92.04 46.10 75.02 zh base 61.19 58.57 64.43 93.88 69.52 50.16 43.42 43.12 86.06 55.69 XLM-R 64.55 55.60 63.98 94.35 69.62 55.42 44.52 44.16 87.73 57.96
7.1 Preliminaries
Inspired by a long debate on the relation between the judgment task and linguistic knowledge of human language users Schütze (2016), we adopt edge probing Tenney et al. (2019b, a) to investigate how much encoders encode syntactic and semantic information, as a potential assessment of linguistic knowledge learned by a machine fine-tuned on the judgment task.
Edge probing focuses on structural labeling tasks in form of span labeling. Given one or two spans, the probing classifier is trained to predict the label with span representations encoded by pre-trained encoders (XLM-R in our case).
For instance, dependency labeling is a typical probing task, but it should be discriminated from dependency parsing. In dependency parsing, the parser should find out: a) the head and dependent, and b) the dependency relation between them. In contrast, in dependency labeling the head and dependent are given so that the task is only about predicting the label between two words. Other tasks follow the same labeling scheme.
7.2 Experiment Settings
We hypothesize that training on MELA can improve the performance of XLM-RoBERTa on syntax-related tasks. To verify this intuition, we design the following experiments.
Probing Tasks
We choose six edge probing tasks: 1) part-of-speech tagging, 2) dependency labeling, 3) constituency labeling, 4) named entity labeling, 5) semantic role labeling, and 6) co-reference.
The data for POS tagging and dependency labeling are multilingual and come from Universal Dependencies V2.13 De Marneffe et al. (2021), including four MELA-high-resource languages (i.e., English, Italian, Russian, and Chinese). The other four tasks are monolingual (in English), and sampled from OntoNotes 5.0 Weischedel et al. (2013).
Experiment 1
We train probing classifiers using span representations from XLM-Rs on English probing tasks. We set the pre-trained XLM-R as control group, and the other four MELA-finetuned XLM-R (trained respectively on four MELA-high-resource languages) as test group.
Experiment 2
For two tasks (pos and dep) with mulitlingual data available, we experiment on cross-lingual transfer. We train probing classifiers on representations from XLM-R in each of four high-resource languages, and run zero-shot evaluation on a target language (lang). We repeat the procedure on XLM-R (see more details in Appendix E.)
7.3 Results
@@
In Experiment 1, we train probing classifiers using representations from different XLM-R variants. The average performance of XLM-R is the lowest across the six edge probing tasks (see Table 5). In Experiment 2, we compare performances of cross-lingual transfer between XLM-R and XLM-R (see Table 6). The results indicate that training on one language of MELA helps the model transfer to that language, and more often than not other languages in part-of-speech tagging and dependency labeling.
@@
MELA-fine-tuned XLM-Rs have better performance on syntax-related probing tasks both in mono-lingual and cros-lingual settings, which supports our hypothesis.
@@ 8 Conclusion
In this work we present MELA, the first multilingual acceptability judgement benchmark covering a diverse set of languages, all annotated by expert linguists. By benchmarking multilingual LLMs on MELA and finetuning XLM-R in different cross-lingual settings, we find that GPT-4 performs on par with supervised XLM-R, and in-language data is crucial, both for few-shot evaluation and supervised finetuning. We probe MELA-finetuned XLM-R for syntax capacity, finding that training on MELA improves the performance on syntax-related probing tasks, which indicates that language models acquire syntactic knowledge during training on linguistic acceptability judgements.
@@
Also, in this work we focused on introducing the MELA dataset and showcasing some of its usages, such as serving as a benchmark for evaluating LLMs and providing a data resource for cross-lingual researches in computational linguistics. We did not propose any new theory, method, or model to improve the understanding of linguistic acceptability in humans or language models. We leave the exploration of other use cases of MELA to future works.
@@
The sentences in MELA, both acceptable and unacceptable, are only intended for researches concerning the acquisition and evaluation of linguistic capabilities (of either humans or language models), and should not be interpreted otherwise. We release MELA under Apache 2.0 license, and note that for the four existing acceptability datasets, RuCoLA is available under Apache 2.0 license, while the authors of CoLA, ItaCoLA, and CoLAC did not provide any license information along with their released datasets. For the individual sentences in MELA, the copyright (where applicable) remains with the original authors or publishers.
Appendix A Prompts
A.1 mTk
The prompt for evaluating mTk is presented in Figure 2, which reuses the prompt for Supernatural Instruction task 616333https://github.com/allenai/natural-instructions/blob/master/tasks/task616_cola_classification.json. In Table 7 we present the results on validation sets when prompting with examples from different languages’ training sets. Each figure is the median of three sets of randomly selected prompts (always one acceptable and one unacceptable).
A.2 Other Open-Source models
Since linguistic acceptability is not included in the finetuning mixture of mT0, BLOOMZ, and Baichuan2, we experiment with several prompts imitating the style of both mTk’s prompt for CoLA and the prompts given by Muennighoff et al. (2023), and the results are presented in Table 13. For the main experiments we use the best prompt selected on this subset of training set, i.e. the seventh prompt in Table 13. For few-shot evaluation, we experiment with both in-language examples and English-only examples, and report the median of three sets of prompts. For the instruction itself we always use English.
A.3 OpenAI Models
For OpenAI models, we use the 0613 version of GPT-3.5-turbo and GPT-4. Due to the limited budget, we choose to use the fifth prompt in Table 13 without further ablations, which is found to perform reasonably well in preliminary experiments. For few-shot evaluation we also experiment with in-language examples and English-only examples. We experimented with different sets of in-context examples with GPT-3.5, and found it to have limited impact: the average MCC for 2-shot evaluation of GPT-3.5 with in-language examples reported in Table 3 is 35.80, and two other runs with different examples yield 37.20 and 36.56, respectively.
prompt en zh it ru de fr es ja ar is avg en 39.13 31.48 12.12 14.92 16.46 12.81 15.77 15.17 6.34 11.21 17.54 zh 18.82 32.18 11.42 8.94 2.96 3.50 17.13 16.92 8.27 13.81 13.39 it 11.67 25.64 18.26 11.54 4.75 8.45 24.30 20.01 9.43 12.62 14.67 ru 14.32 26.47 11.37 11.83 4.39 10.34 15.56 20.60 10.01 16.09 14.10 de 15.39 24.40 12.09 9.05 9.91 6.80 19.59 15.63 7.89 13.40 13.42 fr 13.29 25.41 13.15 12.02 9.29 13.09 17.52 13.42 8.86 13.04 13.91 es 15.09 26.78 14.15 13.83 6.78 6.99 24.42 18.14 14.76 11.64 15.26 ja 13.52 26.74 12.15 2.99 0.59 2.52 11.68 22.45 4.00 12.26 10.89 ar 22.14 25.56 16.81 13.32 9.05 12.61 17.82 16.30 12.72 7.37 15.37 is 9.89 23.25 9.16 5.54 4.79 5.60 13.14 15.53 6.84 15.54 10.93 in-lang. 39.13 32.18 18.26 11.83 9.91 13.09 24.42 22.45 12.72 15.54 19.95
| Model | Examples | en | zh | it | ru | de | fr | es | ja | ar | is | avg |
| mT0 | None | 7.32 | 20.13 | 10.83 | 1.95 | 10.28 | 0.39 | 9.32 | 13.71 | 0.04 | 4.95 | 7.89 |
| In-language | -3.64 | -2.51 | -6.52 | -3.12 | -6.27 | -10.44 | -5.27 | -5.90 | -7.50 | -3.46 | -5.46 | |
| English | -3.64 | -2.51 | -6.52 | -3.12 | -6.27 | -10.44 | -5.27 | -5.90 | -7.50 | -3.46 | -5.46 |
Appendix B Training Details

### v3

@@ 4 Evaluating LLMs with MELA
In this section, we report the performance of fine-tuned XLM-R and several LLMs, open-source or close-source, on MELA.
@@ 4.1 Experimental Settings
To establish a supervised baseline, we use XLM-RoBERTa (Conneau et al., 2020), which is a multilingual version of RoBERTa (Liu et al., 2019) pre-trained on 2.5TB CommonCrawl corpus covering one hundred languages. XLM-R is fine-tuned on the combined training sets of all languages in MELA v1.0 with the hyper-parameters described in Appendix A.2.
For open-source LLMs, we consider BLOOMZ (Scao et al., 2022; Muennighoff et al., 2023), two instruction finetuned variants of mT5 (Xue et al., 2021)—namely mTk (Wang et al., 2022) and mT0 (Muennighoff et al., 2023)—and Baichuan2-Chat (Yang et al., 2023) along with its base model. BLOOMZ is both pre-trained and fine-tuned on 46 languages, which only covers 5 languages in MELA: English, Chinese, French, Spanish, and Arabic444Muennighoff et al. (2023) examine BLOOM’s pre-training corpus ROOTS and estimate it to also contain a small amount of Russian, German, Italian, and Japanese.. The pre-training corpus of mT5 includes all 10 languages in MELA, but mT0 is fine-tuned on the same instruction dataset as BLOOMZ. mTk’s fine-tuning data, on the other hand, covers nine languages in MELA (except for Icelandic) and includes the English CoLA dataset. For Baichuan2, the exact language distribution of pre-training and fine-tuning data is not disclosed. For close-source models, we consider GPT-3.5 and GPT-4o (OpenAI, 2023).
There are several decisions to make when evaluating the above LLMs on MELA: prompt selection and the number of examples in the few-shot scenario. After some pilot experiments, which we describe in Appendix A, we opt to use a binary-choice method with the best performing prompt on the development set, and report the results on the test set in both zero-shot and two-shot scenarios.
4.2 Main results
Results of fine-tuned XLM-R and LLMs evaluated on MELA are given in Table 3. We make the following observations.
Observation 1: GPT-4o exhibits a strong multilingual ability for acceptability judgement.
It achieves the best performance on each individual language in MELA, exceeding supervised fine-tuned XLM-R. Its performance is 11 points higher than finetuned XLM-R even in a 0-shot setting. There is a bigger gap between GPT-4o and finetuned XLM-R on low-resource languages than high-resource ones, likely due to the small amount of training data (500 examples) for XLM-R. Compared to GPT-4o, GPT-3.5 seems to be more English-centric, with drastic performance drop in non-English languages.
Observation 2: LLMs benefit more from in-language examples in two-shot setting.
Our results suggest that prompting with two English examples (most of the time) leads to a lower performance than prompting with in-language examples. On Icelandic, for example, the MCC of the 2-shot setting with English in-context examples is even lower than the 0-shot performance for Baichuan2, GPT-3.5, and GPT-4o.
Observation 3: Baichuan2-Base requires in-context examples
As shown in Table 3, under zero-shot setting, Baichuan2-Base shows random performance555Baichuan2-Base always chooses “B. Unacceptable” as the answer, under all prompts we tested., while its Chat model exhibits non-trivial performance, even for languages such as Spanish. The Base model benefits more from in-context learning examples though, surpassing the Chat model in two-shot settings.
5 Cross-lingual Transfer and Multilingual Fine-tuning
In this section, we investigate cross-lingual transfer and multilingual fine-tuning of linguistic acceptability with XLM-R. All training and evaluation are done on MELA v1.0 as it requires a training set.
5.1 Experimental Settings
Cross-lingual Transfer
To observe the transfer of acceptability judgements across languages, we fine-tune XLM-R on one language, and evaluate on all 10 development sets. We report the median MCC of seven runs for all results to mitigate inter-run variance.666Training details can be found in Appendix A.2.
Multilingual Fine-tuning
We downsample sentences in each language to the same number, and fine-tune XLM-R in three settings: (1) in-language, where the fine-tuning and evaluation languages are the same; (2) all-language, where the model is fine-tuned on a mixture of data containing an equal number of sentences from ten languages; and (3) all-but-in-language, where the model is fine-tuned on a mixture of data containing an equal number of sentences from nine languages, except the one being evaluated on.
@@ 5.2 Results
Results for cross-lingual transfer and multilingual finetuning of XLM-R are reported in Table 4 and Figure 1. We make the following observations.
Observation 1: Cross-lingual transfer is non-trivial.
In Table 4, we see that all numbers are (much) greater than 0, suggesting that transferring from language A to language B is possible, even for acceptability judgment tasks. For instance, fine-tuning XLM-R on 500 Icelandic examples results in 23.16 MCC for a completely unrelated language, Mandarin Chinese. A similar conclusion can be drawn from Figure 1, where the green line, which has no in-language training data, demonstrates an increasing trend for all languages, except for French, which plateaus at around 10 MCC.
Observation 2: Size of training set matters, but not always.
The overall performance when high-resource languages are used as training data (6k training examples), 29.69 MCC, is higher than when low-resource ones are used (500 examples), 18.18 MCC, as shown in the last block of Table 4. However, it must be pointed out that sometimes a (14 times) larger training set does not lead to better performance. For instance, in the second column of Table 4, when evaluated on Chinese, 500 examples of German or Spanish achieve roughly 37 MCC, which is on par with having more than 7,000 training examples for Italian and Russian, with 34.86 and 39.77 MCC respectively. Similarly, when evaluated on Icelandic, 500 German examples again demonstrate a performance on par with many thousands of Chinese, Italian or Russian examples (second to last column of Table 4). Thus the transferring performance between two languages seems to be a result of both the language pair in question as well as the size of the training set.
Observation 3: Among low-resource languages, Arabic training data has the lowest average performance (Table 4).
This is likely due to the fact that Arabic is from a different language family from all other nine languages. From the last column of Table 4, we observe that German and Spanish training data have the best performance, likely because MELA has three Germanic languages and three Romance languages, which may make cross-lingual transfer easier among these cognate languages.
Observation 4: For Italian, Japanese and Arabic, during multi-task training, adding in-language data does not affect performance much.
From Figure 1, we see that the green and orange lines cross for these three languages, suggesting that when training with mixed-language data, in-language examples may not be very critical for these languages.
Task base en it ru zh pos 92.87 0.90 0.60 0.30 1.08 dep 89.41 0.93 0.72 0.51 0.45 const 78.54 0.56 0.10 0.72 0.42 name 93.49 0.74 0.15 1.04 0.59 srl 77.93 4.41 2.07 3.31 2.35 coref 83.84 1.71 0.28 0.14 0.69 avg 86.01 1.54 0.57 1.00 0.93
Probing task Part-of-speech tagging Depedency labeling eval / train en it ru zh avg en it ru zh avg en Mbase 92.87 75.77 65.63 43.33 69.40 89.41 74.99 60.67 40.05 66.28 M 93.77 81.43 68.22 44.66 72.02 90.34 77.40 61.84 45.44 68.76 it Mbase 83.26 94.61 66.90 38.73 70.88 78.17 91.50 60.65 32.35 65.67 M 85.60 95.71 63.73 39.70 71.19 83.56 92.46 62.85 37.31 69.05 ru Mbase 82.97 79.90 95.53 53.18 77.90 77.72 78.86 90.90 42.77 72.56 M 85.42 81.01 95.43 54.06 78.98 80.65 81.27 92.04 46.10 75.02 zh Mbase 61.19 58.57 64.43 93.88 69.52 50.16 43.42 43.12 86.06 55.69 M 64.55 55.60 63.98 94.35 69.62 55.42 44.52 44.16 87.73 57.96
6 Edge Probing
In this section, we adopt edge probing (Tenney et al., 2019a, b) to explore whether fine-tuning on acceptability judgment tasks injects syntax-related information into the pre-trained XLM-R.
@@ 6.1 Experimental settings
Edge probing focuses on structural labeling tasks in the form of span labeling. We choose following tasks: 1) part-of-speech tagging, 2) dependency labeling, 3) constituency labeling, 4) named entity labeling, 5) semantic role labeling, and 6) co-reference.777Tasks 1-2 are from UD (De Marneffe et al., 2021); Tasks 3-6 are from OntoNotes (Weischedel et al., 2013). Take dependency labeling as an example, representations of a dependent and its head, encoded by an XLM-RoBERTa, are used to train a probe classifier to predict the dependency relation between the two words.
We hypothesize that training on MELA can improve the performance of XLM-R on the syntax-related probing tasks above, and design the following experiments.
Experiment 1
We train probing classifiers using span representations from XLM-Rs on English probing tasks. We set the pre-trained Mbase (pre-trained XLM-R) as the control group, and the other four MELA-fine-tuned M as the test group, where specifies the training data of which language from MELA were used for fine-tuning.
Experiment 2
For two tasks (pos and dep) with multilingual data available, we experiment on cross-lingual transfer as well. We train probing classifiers on representations from Mbase in each of four high-resource languages and run zero-shot evaluation on a target language (lang). We repeat the procedure on M (see more details in Appendix B).
@@ 6.2 Results
In Experiment 1 we train probing classifiers using representations from different XLM-R variants, some of which have been fine-tuned on MELA while the base model has not. Results in Table 5 show that the average performance of XLM-Rbase on the six probing tasks is the lowest across the six edge probing tasks (see the last row). We further observe from Table 5 that the semantic role labeling task benefits most from MELA-fine-tuning.
@@
In Experiment 2, the performances of probing tasks are evaluated in the cross-lingual transfer setting. We compare the the pre-trained XLM-R model (Mbase) and XLM-Rs fine-tuned on the linguistic acceptability judgement task of a specific language (M) (see Table 6). The results indicate that the probing classifiers trained on span representations from fine-tuned XLM-R models achieve better performance than the base model. Fine-tuning on one language of MELA helps the model transfer to that language, and more often than not other languages in part-of-speech tagging and dependency labeling.
@@
MELA-fine-tuned XLM-Rs perform better on syntax-related probing tasks in mono-lingual and cross-lingual settings, supporting our hypothesis.
@@ 7 Conclusion
In this work we present MELA, the first multilingual acceptability judgement benchmark covering a diverse set of 10 languages, all annotated by expert linguists. By benchmarking multilingual LLMs on MELA and fine-tuning XLM-R in different cross-lingual settings, we find that (1) GPT-4o ourperforms supervised XLM-R, especially on low-resource languages, that (2) in-language data is crucial for few-shot evaluation and that (3) cross-lingual transfer is non-trivial for all language pairs in supervised fine-tuning. We probe MELA-fine-tuned XLM-R for the syntax information encoded, finding that training on MELA improves the performance on syntax-related probing tasks, which indicates that language models acquire syntactic knowledge during training on linguistic acceptability judgements.
@@
Also, in this work we focused on introducing the MELA dataset and showcasing some of its usages, such as benchmarking LLMs and providing a data resource for cross-lingual research in computational linguistics. We leave the exploration of other use cases of MELA to future work.
@@
The sentences in MELA, both acceptable and unacceptable, are only intended for research concerning the acquisition and evaluation of linguistic capabilities (of either humans or language models), and should not be interpreted otherwise. For individual sentences in MELA, the copyright (where applicable) remains with the original authors or publishers. We ask researchers who use MELA to also cite the original source, i.e., CoLA (Warstadt et al., 2019), ItaCoLA (Trotta et al., 2021), RuCoLA (Mikhailov et al., 2022) and CoLAC (Hu et al., 2023).
Acknowledgments
We thank Bingjie Mao, Jinchi Jiang, Wen Shi, Jialin Guo, Chenhui Liu and Licen Liu for their help in data collection. We also appreciate the suggestions and comments from the anonymous reviewers. This study is supported by Shanghai Pujiang Program awarded to Hai Hu (22PJC063). Ziyin Zhang and Rui Wang are partially supported by the National Natural Science Foundation of China (62176153) and the Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102, as the MoE Key Lab of Artificial Intelligence, AI Institute, Shanghai Jiao Tong University).
Appendix A Benchmark Details
In this section, we provide details about how we make decisions to benchmark LLMs, which includes prompt selection, and the number of examples in the few-shot scenario, along with details of fine-tuning XLM-R.
A.1 Large Language Models
Following MMLU Hendrycks et al. (2021a), we evaluate MELA in the multiple-choice format. We use prompts that end with “Answer:”. Models are required to produce probabilities for index tokens “A” and “B”. The index token with a higher probability is regarded as the decision of models. We tune prompts with open-sourced LLMs on the validation set in our pilot experiment.
We first experiment on 1,000 samples of MELA (50 samples per label per language), using 4 different prompts in 0/2/4/6/8-shot (equal number of positive and negative examples) scenarios with in-language examples provided. The prompt with the highest average MCC is selected to evaluate LLMs on the whole MELA test set (see Figure 2 and 3). The results indicate that 1) prompt-8 is better than others and 2) models no longer improve with more than 2 in-context examples.
Therefore, we carry out formal experiments with prompt-8 (see Figure 4) for both open-sourced LLMs and GPT models in zero and two-shot scenarios (see §4).
prompt model size examples en zh it ru de fr es ja ar is avg ours mTk2 13B in-lang. 22.74 8.47 10.24 16.66 11.96 9.28 13.34 12.00 4.87 10.93 12.05 mTk2 13B en 22.74 8.36 8.98 15.69 14.54 12.30 9.28 10.92 6.52 5.99 11.53 origin mTk2 13B in-lang. 39.13 32.18 18.26 11.83 9.91 13.09 24.42 22.45 12.72 15.54 19.95 mTk2 13B en 39.13 31.48 12.12 14.92 16.46 12.81 15.77 15.17 6.34 11.21 17.54
Note that mTk includes the 2-shot CoLA task in its training data. We also reuse the prompt for Supernatural Instruction task 616888https://github.com/allenai/natural-instructions/blob/master/tasks/task616_cola_classification.json. We compare the results of mTk’s origin CoLA prompt (see Figure 4) and our multiple-choice prompt (see Table 7).
A.2 XLM-R Fine-tuning Details